# Total variation distance between SDEs with stable noise and Brownian motion with applications to Poisson PDEs

### v1

@@ 3. Gradient estimates of $Q_{t}$ and comparison of the generators
(4.3) $\begin{split}\sup_{h\in\mathscr{C}_{b}(1)}&\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\leq Cd^{3/2}\int_{0}^{t}(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\mathrm{d}s\\ &\quad\mbox{}+\frac{Cd}{\alpha-1}(2-\alpha)\int_{0}^{t}s^{-1/2}\,\mathrm{d}s\\ &\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right)+\frac{2Cd}{\alpha-1}(2-\alpha)t^{1/2}\\ &\leq\frac{C_{1}d^{3/2}}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right).\end{split}$
@@
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s,$
@@
(4.4) $\begin{split}\sup_{x\in{}^{d}}&\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\\ &\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sum_{k=1}^{\lfloor t\rfloor-1}\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t-\lfloor t\rfloor}\left\|P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{\lfloor t\rfloor+s}h\right\|_{\infty}\mathrm{d}s,\end{split}$
@@
$\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha).$
@@
$\leq C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}k}\int_{0}^{1}\left((1-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s$
$=C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}k}\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)$
$\leq\frac{3C_{1}d^{3/2}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k},$
@@
$\leq C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\int_{0}^{t-\lfloor t\rfloor}\left((t-\lfloor t\rfloor-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s$
$=C_{1}d^{3/2}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\left((t-\lfloor t\rfloor)^{1-1/\alpha}\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}(t-\lfloor t\rfloor)\right)$
$\leq C_{1}d^{3/2}(2-\alpha)\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)$
$\leq\frac{3C_{1}d^{3/2}}{\alpha-1}(2-\alpha).$
@@
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}$ $\leq\frac{C_{1}d^{3/2}}{\alpha-1}(2-\alpha)\left(1+\sum_{k=1}^{\lfloor t\rfloor-1}\mathrm{e}^{-C_{2}k}\right)$
$\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha).$
@@
$\left\|\mathrm{law}(Y^{x}_{t})-\mathrm{law}(Y^{y}_{t})\right\|_{\mathrm{TV}}\leq 2\wedge\left[C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|\right].$

### v2

@@ 3. Gradient estimates of $Q_{t}$ and comparison of the generators
(4.3) $\begin{split}\sup_{h\in\mathscr{C}_{b}(1)}&\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\leq Cd\log(1+d)\cdot\int_{0}^{t}(t-s)^{-\frac{1}{\alpha}}\,s^{-\frac{1}{2}}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\mathrm{d}s\\ &\quad\mbox{}+\frac{Cd}{\alpha-1}(2-\alpha)\int_{0}^{t}s^{-1/2}\,\mathrm{d}s\\ &\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right)+\frac{2Cd}{\alpha-1}(2-\alpha)t^{1/2}\\ &\leq\frac{C_{1}d\log(1+d)}{\alpha-1}(2-\alpha)t^{-\frac{1}{\alpha}+\frac{1}{2}}\left(1+\log\tfrac{1}{t}\right).\end{split}$
@@
$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s,$
@@
(4.4) $\begin{split}\sup_{x\in\real^{d}}&\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\\ &\leq\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sum_{k=1}^{\lfloor t\rfloor-1}\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\mathrm{d}s\\ &\quad\mbox{}+\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{t-\lfloor t\rfloor}\left\|P_{t-\lfloor t\rfloor-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{\lfloor t\rfloor+s}h\right\|_{\infty}\mathrm{d}s,\end{split}$
@@
$\sup_{h\in\mathscr{C}_{b}(1)}\int_{0}^{1}\left\|P_{1-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h\right\|_{\infty}\mathrm{d}s\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
@@
$\leq C_{1}(2-\alpha)\mathrm{e}^{-C_{2}k}\int_{0}^{1}\left((1-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s\cdot d\log(1+d)$
$=C_{1}(2-\alpha)\mathrm{e}^{-C_{2}k}\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)d\log(1+d)$
$\leq\frac{3C_{1}d\log(1+d)}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k},$
@@
$\leq C_{1}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\int_{0}^{t-\lfloor t\rfloor}\left((t-\lfloor t\rfloor-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{d}s\cdot d\log(1+d)$
$=C_{1}(2-\alpha)\mathrm{e}^{-C_{2}\lfloor t\rfloor}\left((t-\lfloor t\rfloor)^{1-1/\alpha}\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}(t-\lfloor t\rfloor)\right)d\log(1+d)$
$\leq C_{1}(2-\alpha)\left(\frac{\alpha}{\alpha-1}+\frac{1}{\alpha-1}\right)d\log(1+d)$
$\leq\frac{3C_{1}d\log(1+d)}{\alpha-1}(2-\alpha).$
@@
$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}$ $\leq\frac{C_{1}d\log(1+d)}{\alpha-1}(2-\alpha)\left(1+\sum_{k=1}^{\lfloor t\rfloor-1}\mathrm{e}^{-C_{2}k}\right)$
$\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
@@
$\left\|\mathrm{law}(Y^{x}_{t})-\mathrm{law}(Y^{y}_{t})\right\|_{\mathrm{TV}}\leq C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|.$